# FlexiSCD: Flexible Surface Code Deformer for Dynamic Defects

### v2

@@ Abstract
Quantum Error Correction (QEC) codes are vital for protecting quantum computers from noise by encoding quantum information redundantly in physical qubits. However, dynamic defects pose a significant challenge, potentially rendering physical qubits faulty during computation across various quantum platforms. Current defect mitigation strategies focus on surface codes, but often fail to fully restore QEC capability or disrupt logical operations.
In our paper, we introduce FlexiSCD, a code deformation framework that integrates adaptive defect mitigation into surface code workflows. It strategically crafts basic deformation instructions based on gauge transformations, enabling optimized processes tailored to specific defects, restoring QEC capability efficiently with fewer qubit resources. We also design an adaptive code layout for efficient logical operations.
[…]
@@ I Introduction
Figure 1: Comparison of FlexiSCD with existing dynamic defect mitigation methods. (a) Illustration of dynamic defects induced by cosmic rays or error drift. (b) Adaptive Surface Code (ASC) [1] employs fixed deformation operation that leads to bad code distance. (c) Q3DE [2] utilizes fixed-scale enlargement on a fixed layout, resulting in poor resource utilization and communication efficiency. (d) FlexiSCD enables adaptive defect removal, enlargement, and layout optimization, providing a more versatile and efficient response to dynamic defects.
@@
A. Defect Removal excludes defective qubits from syndrome measurements during QEC cycles to ensure the accuracy of decoding and subsequent correction. One leading defect removal technique is ASC (Adaptive Surface Code) [1], which performs a simple transformation (known as super-stabilizer [33, 34, 35, 36, 37, 38]) for defect removal. Defect removal methods such as ASC face two issues:
@@
Prompted by these challenges, we introduce FlexiSCD, a code deformation framework that extends the current surface code instruction set [40, 41, 42, 43] (detailed in Section II-C) to include adaptive defect mitigation capabilities. FlexiSCD meticulously crafts four deformation instructions (Section IV) that provides superior granularity and versatility compared to fixed operations in ASC and Q3DE.
FlexiSCD consists of two components: a compile-time layout generator and a runtime code deformation unit. The code deformation unit uses the extended instruction set to implement a highly optimized deformation strategy that unifies the defect removal and the enlargement approaches. This unified approach eliminates error sources from defects while restoring code distance, overcoming the limitations of ASC and Q3DE (Issue A.(1), B.(1)). Furthermore, due to FlexiSCD’s instruction design, the code deformation unit can effectively manage various defect types and facilitate adaptive enlargement, substantially improving error correction capabilities and resource efficiency (Issue A.(2), B.(2)). Additionally, the layout generator strategically adjusts the spacings between logical qubit patches to ensure unimpeded execution of two-qubit logical operations, thereby maintaining high end-to-end system efficiency (Issue B.(3)). As a result, FlexiSCD resolves all issues of previous methods.
@@
  • We introduce FlexiSCD, a deformation framework designed for adaptive defect mitigation on surface codes.
  • We propose four meticulously crafted deformation instructions , offering highly flexible deformation strategies when combined.
@@ II-A Surface Codes
Impact of Defects. Dynamic defects manifest on either data or syndrome qubits, located within the code or along its boundaries (edges or corners), resulting in intricate defect patterns (Fig. 2(b)). Defective syndrome qubits with high error rates fail to provide accurate error information, which renders the stabilizer inoperable and may compromise the code distance by half or more. Defective data qubits consistently generate excessive errors that may surpass the code’s QEC capability, elevating the logical error rate by orders of magnitude [2].
II-B Gauge Transformation
Gauge transformation [48] is a fundamental technique that underpins FlexiSCD as well as facilitates logical operations on surface codes. In this section, we will introduce gauge operators and basic gauge transformations that can be viewed as atomic instructions on surface codes.
Gauge operators. Gauge refers to “degrees of freedom” in QEC. In surface codes, all Pauli operators that preserve logical qubit states are either stabilizers or logical operators, leaving no spare degrees of freedom. For purposes like defect mitigation, stabilizers can be converted to gauge operators, indicating that they are “set free”. This conversion can be simply done by ceasing the measurement of the selected stabilizers in QEC cycles. Gauge operators form anti-commutative pairs [49], meaning when a stabilizer is converted to a gauge operator, a corresponding Pauli operator that anti-commutes with it (but commutes with the rest of the stabilizers) is also introduced as a gauge operator (or more precisely, a generator of the gauge group [48]). One important utility of gauge operators is to provide freedom to measure stabilizers, $i.e.$, sometimes we can measure two gauge operators of which the product equals a stabilizer for syndrome measurement.
Gauge Transformation explores the degrees of freedom introduced by gauge operators. Converting stabilizers to gauge operators, as mentioned above, is an example of gauge transformations. There are four basic types of gauge transformation and they can be viewed as atomic instructions for performing operations on surface codes:
@@ II-B Gauge Transformation
(1) Stabilizer to Gauge (S2G): S2G converts a stabilizer to a gauge operator. For example, in Fig. 3(c), the stabilizer $s_{1},s_{2},g_{1},g_{2}$ are converted to gauge operators. Meanwhile, $X_{5}$ and $Z_{5}$, which anti-commute with $g_{1},g_{2}$ and $s_{1},s_{2}$, are introduced. The updated stabilizer and gauge set become $S=\{\}$ and $G=\{X_{5},Z_{5},s_{1},s_{2},g_{1},g_{2}\}$. This instruction potentially reduces the code size. S2G instructions are commutative, as they only depend on the set of newly introduced gauges.
(2) Gauge to Stabilizer (G2S): G2S converts a gauge operator into a stabilizer and removes all gauge operators that anti-commute with it. G2S is the inverse process of S2G (Fig. 3(b)(c)). This instruction eliminates redundant degrees of freedom and expands the code.
@@
(4) Gauge to Gauge (G2G): G2G multiplies two gauge operators to form a new gauge operator. For instance, in Fig. 3, the products of gauge operators $s_{1}X_{5}$, $s_{2}X_{5}$, $g_{1}Z_{5}$, and $g_{2}Z_{5}$ also result in gauge operators. This instruction reorganizes gauge operators, potentially separating some qubit from the code. For example, the G2G transformation in Fig. 3(d) indicates that qubit $q_{5}$ can be separated from the code.
In summary, these four atomic instructions offer the building blocks for composing versatile and complex instructions to expand, contract, and deform the surface codes for purposes like performing logical operations and defect mitigation. Viewed as atomic instructions, they can be supported by any surface code implementation by simply altering the syndrome measurement patterns in QEC cycles.
II-C Quantum Program Execution on Surface Codes
@@
Figure 4: (a) Example of merging two code patches. (2) Perform a logical CNOT gate on two distant logical qubits through an ancilla path connecting them. The control (resp. target) logical qubit interact with the ancilla logical qubit on the edge of $Z$- (resp. $X$-) logical operators, with equal-length interacting edges.
Instruction set. Current implementations of surface code logical operations are predominantly based on the Lattice Surgery (LS) method [39], which provides an instruction set to expand and connect surface code patches to update and entangle logical states [48]. These instructions, such as grow, merge, and split, are composed of the atomic instructions introduced in Section II-B. For example, the merge instruction connects two arrays of physical qubit registers through a series of S2G and G2S transformations. In Fig. 4(b), applying merge to the blue and red lines activates syndrome measurements on the ancillary patches between the logical surface code patches, facilitating a logical 2-qubit CNOT operation.
Code layout. Current surface code compilers as well as Q3DE arrange logical qubits on a grid with an inter-space of $d$, as shown in Fig. 4(b). This layout ensures that activated ancillary patches for logical operations maintain the same code distance as logical qubits, providing same level of noise protection.
III The FlexiSCD framework
Figure 5: FlexiSCD integrates a Qubit Layout Generator during compilation time and a Code Deformation Unit during runtime into the current surface code workflow. They are combined to extend adaptive defect mitigation functionality to the existing surface code implementation.
The FlexiSCD framework consists of two main components: a compilation time layout generator and a runtime code deformation unit. As illustrated in Fig. 5, FlexiSCD integrates seamlessly with the existing surface code components, including the algorithm compiler and the execution unit.
@@ III The FlexiSCD framework
The code deformation unit is the core component of FlexiSCD. It deforms logical qubits dynamically to optimize defect mitigation during runtime. Utilizing real-time defect information from defect detectors and the current surface code configuration, the code deformation unit applies our adaptive deformation strategy. This strategy employs instructions from our extended instruction set to effectively remove defects and execute adaptive enlargements. The unit then forwards these instruction schedules to the execution unit for actual updates. Notably, despite the adaptiveness of our deformation methodology, actual code deformation updates can be implemented within a single QEC cycle.
@@
IV The FlexiSCD Instruction Set
Previous methods ASC and Q3DE are limited by their simplistic and fixed deformation operations for diverse defect patterns. In response, FlexiSCD introduces four carefully crafted instructions (Table I) to extend the current surface code instruction set (Section II-C). DataQ_RM and SyndromeQ_RM target the removal of interior data and syndrome qubits, respectively, while PatchQ_RM and PatchQ_ADD modify the boundaries of code patches to achieve adaptive boundary defect removal and code enlargement. Akin to CISC in classical computing, each instruction is constructed from multiple atomic instructions (Section II-B) adapted for the surface code topology, as detailed in Fig. 6. Meanwhile, these instructions are granular enough to ensure seamless integration and versatile enough to support a broad spectrum of deformation operations.
@@ IV The FlexiSCD Instruction Set
ASC DataQ_RM Logical operations, Fixed qubit removal
FlexiSCD DataQ_RM,  SyndromeQ_RM, PatchQ_RM,  PatchQ_ADD Logical operations, Adaptive qubit removal, Adaptive enlargement
IV-A Code Deformation Instructions of FlexiSCD
Fig. 6 presents four deformation instructions introduced by FlexiSCD. Physical qubits are denoted by lattice coordinates, and boundaries by a list of qubit coordinates, both serving as input parameters of instructions.
(a) Data qubit removal (DataQ_RM): DataQ_RM removes a single data qubit. As shown in Fig. 6 (a), it first uses 4 S2G to convert stabilizers $s_{1},s_{2},g_{1},g_{2}$ into gauge operators, while introducing gauge operators $Z_{0},X_{0}$ that anti-commute with $s_{1},s_{2}$ and $g_{1},g_{2}$, respectively. It then uses 4 G2G to adjust gauge operators $s_{1},s_{2},g_{1},g_{2}$ (rectangles) into $X_{0}s_{1},X_{0}s_{2},Z_{0}g_{1},Z_{0}g_{2}$ (triangles), effectively separating $q_{0}$ from the code and enabling its removal. Notably, DataQ_RM coincides with ASC’s super-stabilizer method for single data qubit removal. However, within our framework, DataQ_RM is constructed explicitly and could be combined with other instructions to perform more versatile deformations, as will be shown in Section V.
@@ IV-A Code Deformation Instructions of FlexiSCD
(3) Boundary qubit removal (PatchQ_RM): PatchQ_RM removes defects on the boundary–for example, the syndrome qubit $q_{5}$ or data qubits $q_{1},q_{2}$ in Fig. 6(c). This is achieved by combining several atomic instructions in Section II-B to deform the boundary, thus excluding the defects from the code patch. Fig. 6(c) presents an example where the boundary defective data qubit $q_{1}$ is removed by deforming the boundary from $X_{L2}$ to $X_{L2}^{\prime}$ (specifically, we add and fix $Z_{1}$ as a stabilizer and update the rest of stabilizers and gauge operators accordingly).
(4) New qubits incorporation (PatchQ_ADD): The instruction PatchQ_ADD introduces new qubits beyond the original code. A new data qubit $q$ is initialized in either the $|0\rangle$ or |+ket|$\rangle$| + ⟩ state, equivalent to introducing a new stabilizer $Z_{q}$ or $X_{q}$. New syndrome qubits are then added along with their associated stabilizers, which update the stabilizers and gauge operators through a series of atomic instructions (Section II-B), resulting in a deformed code containing additional qubits. Fig. 6(d) presents an example of adding two data qubits and one syndrome qubit. This instruction can be iterated to facilitate code enlargement.
@@
Figure 6: Examples of the four deformation instructions of FlexiSCD. Qi is the coordinate of qubit $q_{i}$ and XLj is the coordinate list that represents the boundary $X_{Lj}$.
V The Code Deformation Unit
The code deformation unit is the core component of FlexiSCD for adaptive defect mitigation. Before every QEC cycle, the code deformation unit receives the current surface code configuration (including the shape and location of logical qubit patches) from the last round of deformation and current defect information from the dynamic defect detector. Leveraging this information, the code deformation unit executes two subroutines sequentially, as illustrated in Fig. 5. The first subroutine, Defect Removal, removes detected qubits from the current code. The second subroutine Adaptive Enlargement takes the resulting code, calculates the new code distance, and adaptively enlarges the code.
V-A Defect Removal Subroutine

### v3

# Surf-Deformer: Mitigating Dynamic Defects on Surface Code via Adaptive Deformation

@@ Abstract
In this paper, we introduce Surf-Deformer, a code deformation framework that seamlessly integrates adaptive defect mitigation functionality into the current surface code workflow. It crafts several basic deformation instructions based on fundamental gauge transformations, which can be combined to explore a larger design space than previous methods. This enables more optimized deformation processes tailored to specific defect situations, restoring the QEC capability of deformed codes more efficiently with minimal qubit resources. Additionally, we design an adaptive code layout that accommodates our defect mitigation strategy while ensuring efficient execution of logical operations.
Our evaluation shows that Surf-Deformer outperforms previous methods by significantly reducing the end-to-end failure rate of various quantum programs by 35× to 70×, while requiring only about 50% of the qubit resources compared to the previous method to achieve the same level of failure rate. Ablation studies show that Surf-Deformer surpasses previous defect removal methods in preserving QEC capability and facilitates surface code communication by achieving nearly optimal throughput.
@@ I Introduction
Figure 1: Comparison of Surf-Deformer with existing dynamic defect mitigation methods. (a) Illustration of dynamic defects induced by cosmic rays or error drift. (b) Adaptive Surface Code (ASC-S) [1] employs fixed deformation operation that leads to bad code distance. (c) Q3DE [2] utilizes fixed-scale enlargement on a fixed layout, resulting in poor resource utilization and communication efficiency. (d) Surf-Deformer enables adaptive defect removal, enlargement, and layout optimization, providing a more versatile and efficient response to dynamic defects.
@@
A. Defect Removal excludes defective qubits from syndrome measurements during QEC cycles to ensure the accuracy of decoding and subsequent correction. One leading defect removal technique is the state-of-the-art Adaptive Surface Code [1](ASC-S) with the boundary transformation of rotated surface code [33], which performs a simple transformation (known as super-stabilizer [34, 35, 36, 37, 38]) for defect removal. Defect removal methods such as ASC-S face two issues:
@@
Prompted by these challenges, we introduce Surf-Deformer, a code deformation framework that extends the current surface code instruction set [40, 41, 42, 43](detailed in section II-D) to include adaptive defect mitigation capabilities. Surf-Deformer meticulously crafts four deformation instructions (section IV) that provides superior granularity and versatility compared to fixed operations in ASC-S and Q3DE.
Surf-Deformer consists of two components: a compile-time layout generator and a runtime code deformation unit. The code deformation unit uses the extended instruction set to implement a highly optimized deformation strategy that unifies the defect removal and the enlargement approaches. This unified approach eliminates error sources from defects while restoring code distance, overcoming the limitations of ASC-S and Q3DE (Issue A.(1), B.(1)). Furthermore, due to Surf-Deformer’s instruction design, the code deformation unit can effectively manage various defect types and facilitate adaptive enlargement, substantially improving error correction capabilities and resource efficiency (Issue A.(2), B.(2)). Additionally, the layout generator strategically adjusts the spacings between logical qubit patches to ensure unimpeded execution of two-qubit logical operations, thereby maintaining high end-to-end system efficiency (Issue B.(3)). As a result, Surf-Deformer resolves all issues of previous methods.
@@
  • We introduce Surf-Deformer, a deformation framework designed for adaptive defect mitigation on surface codes. [link]
  • We propose four meticulously crafted deformation instructions , offering highly flexible deformation strategies.
@@ II-A Surface Codes
II-B Quantum Errors
Physical Errors. In QEC literature, physical errors typically refer to errors with a low error rate around $10^{-3}\sim 10^{-4}$ and are detected indirectly through stabilizer measurements. Extensive research demonstrates that these physical errors can be effectively corrected by surface codes, resulting in an exponentially suppressed logical error rate as qubit redundancy increases. This applies to various error types, including stochastic errors (single-bit error, CNOT error, etc.)[15] and correlated, non-Markovian errors [48, 49]. Consequently, low-error-rate physical qubits continue to function correctly in the surface code due to QEC protection.
Defects. In contrast, defects are errors with a high error rate or those rendering qubits inoperable. Defects are classified as static defects (e.g., fabrication errors [37, 50], qubit loss [36, 51]) that happen before runtime, and dynamic defects (e.g., cosmic-ray events [22, 21], leakage [25], error drift [27]) that occur during runtime. These defects significantly compromise the functionality of surface codes. For instance, in fig. 2(b), defective syndrome qubits with high error rates fail to provide accurate error information, rendering the stabilizer inoperable and potentially reducing the code distance by half or more. Defective data qubits consistently generate excessive errors that can surpass the code’s QEC capability, elevating the logical error rate by orders of magnitude [2]. Therefore, excluding defects is crucial to maintaining the efficacy of surface codes. Fortunately, these defects can be detected directly using hardware technologies [31, 32] or indirectly through statistical methods [2, 1], enabling alternative strategies beyond standard QEC for addressing them.
II-C Gauge Transformation
This section introduces gauge transformation [52, 53], a fundamental technique that underpins Surf-Deformer. Since our goal is to perform runtime code deformation, it is crucial to preserve the information stored in the logical qubit. To achieve this, we formalize atomic gauge transformation instructions that adjust the stabilizers to be measured while maintaining the logical state encoded in the surface code.
Gauge operators. In classical QEC theory, introducing each stabilizer removes a degree of freedom from the system, allowing the remaining degrees of freedom to encode logical qubits. More generally, introducing fewer stabilizers can save degrees of freedom to encode gauge qubits in addition to logical qubits. These extra degrees of freedom provide more flexibility in code design and can lead to better performance [54, 55]. Operations that change the states of gauge qubits without affecting logical qubits are called gauge operators. Notably, gauge operators form anti-commuting pairs but commute with the other stabilizers [53, 56]. Unlike stabilizers, they are not measured in every QEC cycle but can be measured in different cycles to collectively infer error information.
Gauge Transformation modifies the set of stabilizer and gauge operators to achieve greater flexibility for the surface code. This process can be broken down to four basic types, serving as atomic instructions for code deformation operations:
@@ II-C Gauge Transformation
(1) Stabilizer to Gauge (S2G): S2G converts a stabilizer to a gauge operator. Specifically, we match the stabilizer with an new anti-commuting operator and they both become gauge operators. For example, in fig. 3(b), $X_{5}$ anti-commutes with $g_{1},g_{2}$, so stabilizers $g_{1},g_{2}$ are converted to gauge operators along with $X_{5}$. Likewise, Stabilizer $s_{1},s_{2}$ becomes gauge operators by adding new $Z_{5}$. As shown in in fig. 3(c), updated stabilizer and gauge set become $S=\{\}$ and $G=\{X_{5},Z_{5},s_{1},s_{2},g_{1},g_{2}\}$. This instruction creates degree of freedoms by introducing more gauge operators. Notably, S2G instructions are commutative, as they only depend on the set of newly introduced gauges.
(2) Gauge to Stabilizer (G2S): G2S converts a gauge operator into a stabilizer by measuring it in every cycle and do correction based on the measurement result [57]. Accordingly, all gauge operators that anti-commute with it are removed from the gauge set because it should commute with all gauge operators to be a new stabilizer. G2S is the inverse process of S2G (fig. 3(b)(c)). This instruction eliminates degrees of freedom by introducing more stabilizers.
@@
(4) Gauge to Gauge (G2G): G2G multiplies two gauge operators to form a new gauge operator. For instance, in fig. 3, the products of gauge operators $s_{1}X_{5}$, $s_{2}X_{5}$, $g_{1}Z_{5}$, and $g_{2}Z_{5}$ also result in gauge operators. This instruction reorganizes gauge operators, potentially separating some qubit from the code. For example, the G2G transformation in fig. 3(d) indicates that qubit $q_{5}$ can be separated from the code, as the only remaining gauges operators acting on it are $X_{5}$ and $Z_{5}$.
In summary, these four atomic instructions offer the building blocks for composing versatile and complex instructions to deform the surface codes for defect mitigation. For the proof of how the gauge transformation preserves the logical state, we recommend referring to the appendix section in the arXiv version [58].
II-D Quantum Program Execution on Surface Codes
@@
Figure 4: (a) Example of merging two code patches. (b) Perform a logical CNOT gate on two distant logical qubits through an ancilla path. The control (resp. target) logical qubit interacts with the ancilla logical qubit on the edge of $Z$- (resp. $X$-) logical operators, with equal-length interacting edges.
Instruction set. Current implementations of surface code logical operations are mainly based on the Lattice Surgery (LS) method [39], which provides an instruction set to expand and connect surface code patches to update and entangle logical states [52]. These instructions, such as grow, merge, and split, are composed of the atomic instructions introduced in section II-C. For example, the merge instruction connects two arrays of physical qubit registers through a series of S2G and G2S transformations. In fig. 4(b), applying merge to the blue and red lines activates syndrome measurements on the ancillary patches between the logical surface code patches, facilitating a logical 2-qubit CNOT operation.
Code layout. Current surface code compilers as well as Q3DE arrange logical qubits on a grid with an inter-space of $d$, as shown in fig. 4(b). This layout ensures that activated ancillary patches for logical operations maintain the same code distance as logical qubits, providing same level of noise protection. Please refer to the introductory paper [59] for more details.
Figure 5: Surf-Deformer integrates a Qubit Layout Generator during compilation time and a Code Deformation Unit during runtime into the current surface code workflow. They are combined to extend adaptive defect mitigation functionality to the existing surface code implementation.
III The Surf-Deformer framework
The Surf-Deformer framework consists of two main components: a compilation time layout generator and a runtime code deformation unit. As illustrated in fig. 5, Surf-Deformer integrates seamlessly with the existing surface code components, including the algorithm compiler and the execution unit.
@@ III The Surf-Deformer framework
The code deformation unit is the core component of Surf-Deformer. It deforms logical qubits dynamically to optimize defect mitigation during runtime. Utilizing real-time defect information from defect detectors and the current surface code configuration, the code deformation unit applies our adaptive deformation strategy. This strategy employs instructions from our extended instruction set to effectively remove defects and execute adaptive enlargements. The unit then forwards these instruction schedules to the execution unit for actual updates. Notably, despite the adaptiveness of our deformation methodology, actual code deformation updates can be implemented within a single QEC cycle.
@@
IV The Surf-Deformer Instruction Set
Previous methods ASC-S and Q3DE are limited by their simplistic and fixed deformation operations for diverse defect patterns. In response, Surf-Deformer introduces four carefully crafted instructions (table I) to extend the current surface code instruction set (section II-D). DataQ_RM and SyndromeQ_RM target the removal of interior data and syndrome qubits, respectively, while PatchQ_RM and PatchQ_ADD modify the boundaries of code patches to achieve adaptive boundary defect removal and code enlargement. Akin to CISC in classical computing, each instruction is constructed from multiple atomic instructions (section II-C) adapted for the surface code topology, as detailed in fig. 6. Meanwhile, these instructions are granular enough to ensure seamless integration and versatile enough to support a broad spectrum of deformation operations.
@@ IV The Surf-Deformer Instruction Set
ASC-S DataQ_RM Logical operations, Fixed qubit removal
Surf-Deformer DataQ_RM,  SyndromeQ_RM, PatchQ_RM,  PatchQ_ADD Logical operations, Adaptive qubit removal, Adaptive enlargement
IV-A Code Deformation Instructions of Surf-Deformer
fig. 6 presents four deformation instructions introduced by Surf-Deformer. Physical qubits are denoted by lattice coordinates, and boundaries by a list of qubit coordinates, both serving as input parameters of instructions.
(a) Data qubit removal (DataQ_RM): DataQ_RM removes a single data qubit. As shown in fig. 6(a), it first uses four S2G to convert stabilizers $s_{1},s_{2},g_{1},g_{2}$ into gauge operators, while introducing gauge operators $Z_{0},X_{0}$ that anti-commute with $s_{1},s_{2}$ and $g_{1},g_{2}$, respectively. It then uses four G2G to adjust gauge operators $s_{1},s_{2},g_{1},g_{2}$ (rectangles) into $X_{0}s_{1},X_{0}s_{2},Z_{0}g_{1},Z_{0}g_{2}$ (triangles), effectively separating $q_{0}$ from the code and enabling its removal. Notably, DataQ_RM coincides with ASC-S’s super-stabilizer method for single data qubit removal. However, within our framework, DataQ_RM is constructed explicitly and could be combined with other instructions to perform more versatile deformations, as will be shown in section V.
@@ IV-A Code Deformation Instructions of Surf-Deformer
Figure 6: Examples of the four deformation instructions of Surf-Deformer. Qi is the coordinate of qubit $q_{i}$ and XLj is the coordinate that represents the boundary $X_{Lj}$.
(3) Boundary qubit removal (PatchQ_RM): PatchQ_RM removes defects on the boundary–e.g. the syndrome qubit $q_{5}$ or data qubits $q_{1},q_{2}$ in fig. 6(c). This is achieved by combining several atomic instructions in section II-C to deform the boundary, thus excluding the defects from the code patch. fig. 6(c) presents an example where the boundary defective data qubit $q_{1}$ is removed by deforming the boundary from $X_{L2}$ to $X_{L2}^{\prime}$ (specifically, we add and fix $Z_{1}$ as a stabilizer and update the rest of stabilizers and gauge operators accordingly).
(4) New qubits incorporation (PatchQ_ADD): The instruction PatchQ_ADD introduces new qubits beyond the original code. A new data qubit $q$ is initialized in either the $|0\rangle$ or |+ket|$\rangle$| + ⟩ state, equivalent to introducing a new stabilizer $Z_{q}$ or $X_{q}$. New syndrome qubits are then added along with their associated stabilizers, which update the stabilizers and gauge operators through a series of atomic instructions (section II-C), resulting in a deformed code containing additional qubits. fig. 6(d) presents an example of adding two data qubits and one syndrome qubit. This instruction can be iterated to facilitate code enlargement.